### v2

@@ 4.3 Deferral using Ensemble Agreement
Ensemble agreement is an instantiation of score based deferral rules Jitkrittum et al. (2023). In its simplest form, we associate a score to the prediction output of an ensemble model and interpret this score as a measure of confidence in that particular prediction. We can then defer to a larger model when the confidence is below a predetermined threshold. For a model $H_{1}^{k}(x)$ and a scoring function $s(x)$, the deferral rule is defined as
@@
where $\mathbb{I}$ is the indicator function. In general, such score based deferral rules can sometimes be misleading. For instance, in multi-class classification, rules that map the outputs of a classifier to a probability distribution over labels can produce confidently incorrect predictions when an unperceivable amount of perturbation is added to the data sample Szegedy (2013). However, such rules have been shown to be effective in practice Wang et al. (2018a); Gangrade et al. (2021); Mamou et al. (2022); Gupta et al. (2024) implying that such adversarial data is rare in many cases. Motivated by this observation, we propose the following property of scoring functions (and by extension the corresponding deferral rule).
@@
We define safe deferral with respect to classification tasks for simplicity. The definition can be extended to other cases with appropriate choice of loss function. Intuitively, this formalizes the notion that the deferral rule can probabilistically identify a subset of $\mathcal{X}\times\mathcal{Y}$ for which $H^{k}_{1}(x)$ is correct. Alternatively, we can think of the deferral rule as a one sided classifier, similar to the problem studied in Goyal et al. (2020), where the class to be identified is the subset of $\mathcal{X}\times\mathcal{Y}$ where $H^{k}_{1}(x)$ is correct. In general, such a safe deferral rule need not exist for a given model $H_{1}^{k}(x)$. However, in Section 5 we demonstrate empirically that such rules can in-fact be constructed without additional training for many real world tasks with appropriate choice of $\theta$ (Figure 7). These rules lead to selection rates as high as 90% of the data in cases such as ImageNet1k.
Assuming access to a safe deferral rule, observe that for $\xi\geq\epsilon>0$, such rules are feasible for the program in Equation 2 (see Proposition 4.1). This implies that every cascade, =$\mathcal{M}={\{H_{1}^{k},h_{2}\}}$ using a safe deferral rule $r$, is competitive with the larger model $h_{2}$ in terms of accuracy. We empirically evaluate these drop-in cascades in Section 5. Note that the optimal safe deferral rule can lead to an improvement in the accuracy of ABC in theory (see, Appendix A), and we see accuracy improvements in our experiments (Section 5).
@@
We evaluate two flavors of deferral rules that capture the notion of agreement between models. In cases where we have direct access to the models, we directly use the outputs produced by models in the ensemble. However, in certain cases, for instance when interfacing through an inference API from a third party provider, we only have black box access to models. For such cases, we use a voting scheme between the models in an ensemble to construct our scoring function. Concretely, for an ensemble $H^{k}$ consisting of $k\geq 1$ models, let $s(x;H^{k})$ denote the average score of the majority prediction on $x$ and let $\text{vote}(x;H^{k})=\frac{1}{k}\sum_{h\in H^{k}}\mathbb{I}[H^{k}(x)=h(x)]$ denote the fraction of votes received by the prediction, where,
@@ 4.4 Inference Cost Savings and Competitiveness
    The ABC classifier is competitive with the large classifier $h_{2}$ in terms of accuracy,
@@
Thus the cost savings we can expect with a drop-in cascade depends on three key factors; the relative cost $\gamma$, the degree of parallelization $\rho$ and the deferral rate $\mathbb{P}(r(x)=1)$, or equivalently, the selection rate $\mathbb{P}(r(x)=0)$. In particular, for every feasible deferral rule $r$, in the best case scenario where the cost of the smaller model is negligible, ($\gamma=0$), the cost of inference reduces by the selection rate, $P(r(x)=0)$. Conversely, in the worst-case scenarios, the cost can be $(k+1)$ times the cost of the larger model. In the next section, we demonstrate real world scenarios where the favorable interplay of these three quantities lead to significant improvements in inference cost. See Appendix C and Figure 7 for information on selection rates for various models and datasets considered here.
@@ 5 Experiments
This section evaluates our training-free Agreement-Based Cascading (ABC) approach across a variety of language and vision tasks, focusing on accuracy and inference cost. First, in Section 5.1.1, we examine ABC’s accuracy-cost tradeoff against state-of-the-art models under full parallelization, setting the parallelization factor $(\rho)$ to $\rho=1.0$ for optimal accuracy-FLOPs trade-off. Next, in Section 5.1.2, we show that using ensembles at lower tiers has minimal impact on inference cost when lower-tier models are at least 50×50$\times$50 × cheaper than the largest model. In such cases, the relative cost $(\gamma)$ satisfies $\gamma\leq\frac{1}{50}$, making ABC effective without parallelization. Finally, Section 5.2 explores practical scenarios where low relative costs make ABC is naturally suited for real-world deployments.
@@
ABC’s deferral rule uses a configurable voting threshold, $\theta$ (see Equations 3 and 4) at each cascading tier. We estimate $\theta$ empirically on a small set of unseen data; see Appendix B for details.
@@
To evaluate ABC, we use a range of benchmark datasets for image and language tasks, as shown in Table 2 in the Appendix. Additional datasets are used in Section 5.2.3 to align with those explored by state-of-the-art baselines.
@@
We select diverse models for both image and language tasks, summarized in Appendix’s Table 3. For BERT-based models, we use the BASE and LARGE, and for image models, we tier by FLOPs count. All models are sourced from HuggingFace Zoo for inference without any additional training effort on our end. Section 5.2.3 uses models from LLaMA 3, Gemma 2, and Qwen 2 families via the Together API. We detail this section’s evaluation setup in Appendix D.2.
@@ 5.1.1 When Parallelization is Cheap
The accuracy vs. FLOPs for ABC is shown in Figure 2. As a point of comparison, we also include Wisdom-of-Committees (WoC) Wang et al. (2021), a popular and representative confidence-based model cascading method. In most cases, we observe that ABC is able to improve the Pareto frontiers, as it usually sees a 11112222 point increase in accuracy. We attribute this to a combination of (a) the improvement ensembles can have on accuracy that is widely noted in literature Gontijo-Lopes et al. (2022); Jiang et al. (2023) and (2) the improvement agreement-based rules can cause in cascading (Appendix A). Overall, in terms of accuracy, ABC either exceeds—or at least matches—the accuracy of the best models at a fixed FLOPs budget, and is a practical drop-in replacement for these models.
@@ 5.1.2 Disparity in Relative Cost is High
Of course, in many real-world scenarios, the cost of evaluating ensembles is not negligible. For instance, in the edge-to-cloud inference scenario discussed in Section 5.2.1, models in an ensemble often are evaluated sequentially. However, even in such scenarios, if the relative cost of models across each level of the cascade, $\gamma$, is small enough, this additional cost becomes negligible compared to the overall cost.
@@
Figure 3: Fraction of inference cost saved as a function of relative cost of models $(\gamma)$, assuming a fixed selection rate $\mathbb{P}(r(x)=0)$. As parallelization decreases from fully parallel $(\rho=1)$ to sequential $(\rho=0)$, cost of evaluating ensembles increase and the cost savings decrease. When models across tiers are of similar size (e.g. smaller model is at most 5×5$\times$5 × smaller, $\gamma\geq\frac{1}{5}$), some parallelization is needed for ABC to reduce costs effectively. However, for lower relative costs (e.g., smaller model is at least 50×50$\times$50 × smaller, $\gamma\leq\frac{1}{50}$), sequential and parallel settings achieve meaningful savings, showing ABC’s efficiency even with the added cost of using ensembles.
@@ 5.2 Real-world Use-cases
As noted in the previous two subsections, ABC  either improves on, or is at least competitive with, the single best model in terms of accuracy. Moreover, whenever the relative cost is small enough, ABC  suffers negligible additional penalty for using an ensemble of models. As noted in Figure 1, accuracy vs inference cost scaling for ML models already imply that the relative cost of models that only differ a few points in accuracy is small. For instance, the state-of-the-art model performance on ImageNet1K dataset attains about $83\%$ top-1 accuracy with $70$B FLOPs. A Pareto-optimal model that achieves $63\%$ accuracy requires only about $1$B flops, and thus a two-level ABC with ensembles of these two models have relative cost of $\gamma=\frac{1}{70}$.
In many real world scenarios, this disparity in relative cost is further amplified due to deployment considerations — for instance, in an edge-to-cloud inference setting, local inter-process communication (IPC) latency is typically two orders smaller than remote (cloud) IPC latency $(\gamma\approx 10^{-2})$. Moreover, the ability of models in ABC to be placed on multiple, distributed devices with negligible synchronization overhead adds more opportunities for inference cost reduction. We consider three such scenarios, their respective cost models, and the benefits of ABC in these cases here. These include edge-to-cloud inference (Section 5.2.1), cloud-based model serving on heterogeneous GPUs (Section 5.2.2), and black-box inference access to model API services (Section 5.2.3).
@@ 5.2.1 Communication Cost in Edge-to-Cloud Inference
An advantage of ABC is that it allows a single large model to be split into multiple, potentially much smaller models, with only a simple reduce operation required to compute agreement. This allows us to tune device placement at various levels of ABC to improve inference costs. One use-case where this is beneficial is edge-to-cloud inference Forooghifar et al. (2019); here, inference requests are generated on user-facing edge devices like mobile phones or smart devices, which are sent over the network to a cloud service for evaluation. Given an inference request generated by a user interaction, the time-to-response or response latency in such case is dominated by communication overheads (network speed, serialization overhead, network congestion, etc.) beyond our control. By using ABC for such applications, we are able to distribute the inference load between tiny, on-device models and the cloud models, allowing us to avoid communication costs for a significant portion of requests.
@@
Our results, as shown in Figure 4, show that the flexibility that ABC affords in terms of model placement allows significant latency reductions, while providing superior accuracy compared to a single cloud model. In particular, we see that cascading in these scenarios provides an  14$\times$ reduction in communication cost for language tasks like SST-2; and for image datasets, we see a 5×5$\times$5 × reduction for ImageNet1K and a 8×8$\times$8 × reduction in CIFAR10.
@@ 5.2.2 Monetary Cost for Model Serving on Heterogeneous Hardware
Another use-case which can take advantage of the model placement flexibility of ABC is using heterogeneous hardware for model serving on the cloud Crago & Walters (2015); Li et al. (2021a; 2023); Mo et al. (2023). GPU/Accelerator hardware typically has a disproportionately large difference in hardware costs compared to their throughput difference. For instance, based on the current pricing model offered by Lambda Lambda (2024), a popular cloud rental platform, the rental pricing of a single A100 is \$1.40/hour and a V100 node is \$0.06/hour ($\gamma\approx 4\times 10^{-2}$), while the rated 32-bit tensor core throughput is 312 TFLOPs for A100 and 125 TFLOPS for V100. In this scenario, a simple placement strategy for a 2-level ABC that reduces inference cost may place the smaller model on V100 nodes and larger models on A100 nodes. Since the lower level models are—at least—an order of magnitude cheaper than the larger one in terms of FLOPs, this offsets the throughput loss we incur when switching from A100 to V100; and as a result, the ABC implementation incurs a much lower average cost.
To concretely evaluate ABC under such a cost model, we retrieve the costs of GPU usage by hour from Lambda Cloud’s pricing (see, Table 4 in the appendix for details). For simplicity, we assume that each ensemble tier is set up on a distinct GPU in increasing order of GPU sophistication, and serves a uniform inference request rate. We also assume that the nodes are co-located and communication cost between them is negligible. We show a summary in Figure 4 and present the detailed results in Table 5, in the Appendix. We can see at least a 3×3$\times$3 × reduction in inference cost in terms of \$/\\$/\$ /hour for image tasks, and a more moderate $10$-$30\%$ reduction in language tasks.
@@ 5.2.3 Monetary Cost for Black-box API-Based Inference
Finally, in the current landscape dominated by LLMs, many providers offer black-box API access to their proprietary LLMs Abdalla et al. (2023); Sun et al. (2022); Hadi et al. (2024). Similarly to the large cost difference between GPU generations discussed in the previous subsection, we also observe a large cost disparity between various generations/tiers of API calls. For example, using Together.ai—one of the lowest-cost serverless endpoint providers for LLMs111together.ai/pricing as of September 2024.—as a case study, we find that models within the 7B-8B range cost \$0.20 per million tokens, while LlaMA3.1-405B costs \$5.00. This translates to the larger model costing 25$\times$ more than the smaller model range $(\gamma=\frac{1}{25},\rho=1)$. If we consider GPT-4 as the gold standard, the cost of usage quickly scales to 150$\times$ of our reference smaller model’s costs. [^2] GPT-4-1106-preview costs \$30 as of September 2024..
@@
Since we only have black-box access to these models, we cannot employ a score-based deferral rule. However, we demonstrate that using ABC’s voting-based rule defined in Section 4 is also effective in such scenarios. For baseline comparison, we use FrugalGPT, 2 variants of AutoMix, and MoT LLM Cascade. These state-of-the-art cascading methods are specifically designed for scenarios where we make API calls to black-box model endpoints, in contrast to our more generally applicable ABC method. We access the models from Together.ai—described in Table 1—for these experiments, and consider a setting that is advantageous to the baselines, by selecting the best singular model from each performance tier in their respective approaches.
All the baseline methods considered here are significantly more complex and involved to set up than ABC; both AutoMix and FrugalGPT involve training a router or deferral rule at each cascade level which has to be repeated for every new task or model change. Unlike AutoMix, FrugalGPT requires training a DistilBERT-based scorer, which would require the user’s possession of GPU resources. MoT LLM Cascade generates multiple results and varies the randomness in the LLM’s responses via sampling, while using in-context demonstrations and reasoning techniques (e.g., Chain-of-Thought Wei et al. (2022)) to influence how the model generates answers. In contrast, ABC uses a much simpler voting-based safe deferral rule, without involving additional training or complex routing strategies. Further method-specific details for these experiments as well as additional results can be found in Appendix D.2.

### v3

@@ 4.3 Deferral using Ensemble Agreement
Ensemble agreement is an instantiation of score based deferral rules (Jitkrittum et al., 2023). In its simplest form, we associate a score to the prediction output of an ensemble model and interpret this score as a measure of confidence in that particular prediction. We can then defer to a larger model when the confidence is below a predetermined threshold. For a model $H_{1}^{k}(x)$ and a scoring function $s(x)$, the deferral rule is defined as
@@
where $\mathbb{I}$ is the indicator function. In general, such score based deferral rules can sometimes be misleading. For instance, in multi-class classification, rules that map the outputs of a classifier to a probability distribution over labels can produce confidently incorrect predictions when an unperceivable amount of perturbation is added to the data sample (Szegedy, 2013). However, such rules have been shown to be effective in practice (Wang et al., 2018a; Gangrade et al., 2021; Mamou et al., 2022; Gupta et al., 2024) implying that such adversarial data is rare in many cases. Motivated by this observation, we propose the following property of scoring functions (and by extension the corresponding deferral rule).
@@
We define safe deferral with respect to classification tasks for simplicity. The definition can be extended to other cases with appropriate choice of loss function. Intuitively, this formalizes the notion that the deferral rule can probabilistically identify a subset of $\mathcal{X}\times\mathcal{Y}$ for which $H^{k}_{1}(x)$ is correct. Alternatively, we can think of the deferral rule as a one sided classifier, similar to the problem studied in Goyal et al. (2020), where the class to be identified is the subset of $\mathcal{X}\times\mathcal{Y}$ where $H^{k}_{1}(x)$ is correct. In general, such a safe deferral rule need not exist for a given model $H_{1}^{k}(x)$. However, in §5 we demonstrate empirically that such rules can in-fact be constructed without additional training for many real world tasks with appropriate choice of $\theta$ (Figure 7). These rules lead to selection rates as high as 90% of the data in cases such as ImageNet-1K.
Assuming access to a safe deferral rule, observe that for $\xi\geq\epsilon>0$, such rules are feasible for the program in Equation 2 (see Proposition 4.1). This implies that every cascade, =$\mathcal{M}={\{H_{1}^{k},h_{2}\}}$ using a safe deferral rule $r$, is competitive with the larger model $h_{2}$ in terms of accuracy. We empirically evaluate these drop-in cascades in §5. Note that the optimal safe deferral rule can lead to an improvement in the accuracy of ABC in theory (see, Appendix A), and we see accuracy improvements in our experiments (§5).
@@
We evaluate two flavors of deferral rules that capture the notion of agreement between models. In cases where we have direct access to the models, we directly use the outputs produced by models in the ensemble. However, in certain cases, for instance when interfacing through a third-party provider’s inference API, we only have black box access to models. For such cases, we use a voting scheme between the models in an ensemble to construct our scoring function. Concretely, for an ensemble $H^{k}$ consisting of $k\geq 1$ models, let $s(x;H^{k})$ denote the average score of the majority prediction on $x$ and let $\text{vote}(x;H^{k})=\frac{1}{k}\sum_{h\in H^{k}}\mathbb{I}[H^{k}(x)=h(x)]$ denote the fraction of votes received by the prediction, where,
@@ 4.4 Inference Cost Savings and Competitiveness
    The ABC classifier is competitive with the large classifier $h_{2}$ in terms of accuracy (zero-one loss),
@@
[…]
@@ 5 Experiments
This section evaluates our training-free Agreement-Based Cascading (ABC) approach across a variety of language and vision tasks, focusing on accuracy and inference cost. First, in §5.1.1, we examine ABC’s accuracy-cost tradeoff against state-of-the-art models under full parallelization, setting the parallelization factor $(\rho)$ to $\rho=1.0$ for optimal tradeoff. Next, in §5.1.2, we show that using ensembles at lower tiers has minimal impact on inference cost when lower-tier models are at least 50×50$\times$50 × cheaper than the largest model. In such cases, the relative cost $(\gamma)$ satisfies $\gamma\leq\frac{1}{50}$, making ABC effective without parallelization. Finally, §5.2 explores practical scenarios where low relative costs make ABC is naturally suited for real-world deployments.
@@
ABC’s deferral rule uses a configurable voting threshold, $\theta$ (see Equations 3 and 4) at each cascading tier. We estimate $\theta$ empirically on a small set of unseen data; see App. B for details.
@@
To evaluate ABC, we use a range of benchmark datasets for image and language tasks, as shown in Table 2 in the Appendix. Additional datasets are used in §5.2.3 to align with those explored by state-of-the-art baselines.
@@
We select diverse models for both image and language tasks, summarized in Appendix’s Table 3. For BERT-based models, we use the BASE and LARGE, and for image models, we tier by FLOPs count. All models are sourced from HuggingFace Zoo for inference without any additional training effort on our end. §5.2.3 uses models from LLaMA 3, Gemma 2, and Qwen 2 families via the Together API. We detail this section’s experimental setup in App. D.2.
Evaluation:
For the generation tasks we consider in §5.2.3, the datasets each consist of a fixed set of possible outputs akin to the classification tasks, and we apply our deferral rule on the final output at each tier. Certain cases don’t have a fixed set of output labels (while still not being open-ended generation), like (1) GSM8K, where there is a final numeric answer at the end of the generated output to the math word problems, and (2) CoQA, where we used F1-score to capture overlaps between predictions and ground-truth answers. We do not discuss open-ended generation tasks in this work as the deferral rule we consider—and the baseline methods that we compare to—is not directly applicable to such cases.
@@ 5.1.1 When Parallelization is Cheap
Figure 3: Fraction of inference cost saved as a function of relative cost of models $(\gamma)$, assuming a fixed selection rate $\mathbb{P}(r(x)=0)$. As parallelization decreases from fully parallel $(\rho=1)$ to sequential $(\rho=0)$, cost of evaluating ensembles increase and the cost savings decrease. When models across tiers are of similar size (e.g., smaller model is at most 5×5$\times$5 × smaller, $\gamma\geq\frac{1}{5}$), some parallelization is needed for ABC to reduce costs effectively. However, for lower relative costs (e.g., smaller model is at least 50×50$\times$50 × smaller, $\gamma\leq\frac{1}{50}$), sequential and parallel settings achieve meaningful savings, showing ABC ’s efficiency even with the added cost of using ensembles.
The accuracy vs. FLOPs for ABC is shown in Figure 2. As a point of comparison, we also include Wisdom-of-Committees (WoC) (Wang et al., 2021), a popular and representative confidence-based model cascading method. In most cases, we observe that ABC is able to improve the Pareto frontiers, as it usually sees a 111222 point increase in accuracy. We attribute this to a combination of (a) the improvement ensembles can have on accuracy that is widely noted in literature (Gontijo-Lopes et al., 2022; Jiang et al., 2023) and (2) the improvement agreement-based rules can cause in cascading (Appendix A). Overall, in terms of accuracy, ABC either exceeds—or at least matches—the accuracy of the best models at a fixed FLOPs budget, and is a practical drop-in replacement for these models.
@@ 5.1.2 Disparity in Relative Cost is High
Of course, in many real-world scenarios, the cost of evaluating ensembles is not negligible. For instance, in the edge-to-cloud inference scenario discussed in §5.2.1, models in an ensemble often are evaluated sequentially. However, even in such scenarios, if the relative cost of models across each level of the cascade, $\gamma$, is small enough, this additional cost becomes negligible compared to the overall cost.
@@
Takeaway #1: Although ensembling requires additional inference costs, these costs can be mitigated when (1) models can be parallelized or (2) smaller models are several magnitudes cheaper than larger ones. In such scenarios, ABC can achieve substantial accuracy improvements while reducing inference costs.
@@ 5.2 Real-world Use-cases
[…]
In many real world scenarios, this disparity in relative cost is further amplified due to deployment considerations — for instance, in an edge-to-cloud inference setting, local inter-process communication (IPC) latency is typically two orders smaller than remote (cloud) IPC latency $(\gamma\approx 10^{-2})$. Moreover, the ability of models in ABC to be placed on multiple, distributed devices with negligible synchronization overhead adds more opportunities for inference cost reduction. We consider three such scenarios, their respective cost models, and the benefits of ABC in these cases here. These include edge-to-cloud inference (§5.2.1), cloud-based model serving on heterogeneous GPUs (§5.2.2), and black-box inference access to model API services (§5.2.3).
@@ 5.2.1 Communication Cost in Edge-to-Cloud Inference
An advantage of ABC is that it allows a single large model to be split into multiple, potentially much smaller models, with only a simple reduce operation required to compute agreement. This allows us to tune device placement at various levels of ABC to improve inference costs. One use-case where this is beneficial is edge-to-cloud inference (Forooghifar et al., 2019); here, inference requests are generated on user-facing edge devices like mobile phones or smart devices, which are sent over the network to a cloud service for evaluation. Given an inference request generated by a user interaction, the time-to-response or response latency in such case is dominated by communication overheads (network speed, serialization overhead, network congestion, etc.) beyond our control. By using ABC for such applications, we are able to distribute the inference load between tiny, on-device models and the cloud models, allowing us to avoid communication costs for a significant portion of requests.
@@
Our results, as shown in Figure 4, show that the flexibility that ABC affords in terms of model placement allows significant latency reductions, while providing superior accuracy compared to a single cloud model. In particular, we see that cascading in these scenarios provides an  14$\times$ reduction in communication cost for language tasks like SST-2; and for image datasets, we see a 5×5$\times$5 × reduction for ImageNet-1K and a 8×8$\times$8 × reduction in CIFAR10.
Takeaway #2: ABC enables model placement flexibility where small models run locally and large models in the cloud. Communication delays can make relative costs significant, resulting in ABC achieving 5-14$\times$ reductions in communication costs while maintaining superior accuracy.
@@ 5.2.2 Monetary Cost for Model Serving on Heterogeneous Hardware
Another use-case which can take advantage of the model placement flexibility of ABC is using heterogeneous hardware for model serving on the cloud (Crago & Walters, 2015; Li et al., 2021a; 2023; Mo et al., 2023). GPU/Accelerator hardware typically has a disproportionately large difference in hardware costs compared to their throughput difference. For instance, based on the current pricing model offered by Lambda (Lambda, 2024), a popular cloud rental platform, the rental pricing of a single A100 is \$1.40/hour and a V100 node is \$0.06/hour ($\gamma\approx 4\times 10^{-2}$), while the rated 32-bit tensor core throughput is 312 TFLOPs for A100 and 125 TFLOPS for V100. In this scenario, a simple placement strategy for a 2-level ABC that reduces inference cost may place the smaller model on V100 nodes and larger models on A100 nodes. Since the lower level models are—at least—an order of magnitude cheaper than the larger one in terms of FLOPs, this offsets the throughput loss we incur when switching from A100 to V100; and as a result, the ABC implementation incurs a much lower average cost.
[…]
Takeaway #3: GPU rental cost differences significantly exceed throughput differences. ABC optimally places models across hardware tiers, achieving 3$\times$ cost reductions for image tasks and 10-30% savings for text tasks.
@@ 5.2.3 Monetary Cost for Black-box API-Based Inference
Finally, in the current landscape dominated by LLMs, many providers offer black-box API access to their proprietary LLMs (Abdalla et al., 2023; Sun et al., 2022; Hadi et al., 2024). Similarly to the large cost difference between GPU generations discussed in the previous subsection, we also observe a large cost disparity between various generations/tiers of API calls. For example, using Together.ai—one of the lowest-cost serverless endpoint providers for LLMs111together.ai/pricing as of September 2024.—as a case study, we find that models within the 7B-8B range cost \$0.20 per million tokens, while LlaMA3.1-405B costs \$5.00. This translates to the larger model costing 25$\times$ more than the smaller model range $(\gamma=\frac{1}{25},\rho=1)$. If we consider GPT-4 as the gold standard, the cost of usage quickly scales to 150$\times$ of our reference smaller model’s costs. [^2] GPT-4-1106-preview costs \$30 as of September 2024..
Since we only have black-box access to these models, we cannot employ a score-based deferral rule. However, we demonstrate that using ABC’s voting-based rule defined in §4 is also effective in such scenarios. For baseline comparison, we use FrugalGPT, 2 variants of AutoMix, and MoT LLM Cascade. These state-of-the-art cascading methods are specifically designed for scenarios where we make API calls to black-box model endpoints, in contrast to our more generally applicable ABC method. We access the models from Together.ai—described in Table 1—for these experiments, and consider a setting that is advantageous to the baselines, by selecting the best singular model from each performance tier in their respective approaches.
All the baseline methods considered here are significantly more complex and involved to set up than ABC; both AutoMix and FrugalGPT involve training a router or deferral rule at each cascade level which has to be repeated for every new task or model change. Unlike AutoMix, FrugalGPT requires training a DistilBERT-based scorer, which would require the user’s possession of GPU resources. MoT LLM Cascade generates multiple results and varies the randomness in the LLM’s responses via sampling, while using in-context demonstrations and reasoning techniques (e.g., Chain-of-Thought (Wei et al., 2022)) to influence how the model generates answers. In contrast, ABC uses a much simpler voting-based safe deferral rule, without involving additional training or complex routing strategies. Further method-specific details for these experiments as well as additional results can be found in Appendix D.2.
@@
Takeaway #4: LLM API pricing creates extreme cost ratios. ABC’s simple voting mechanism achieves 2-25× cost reductions compared to SOTA cascading methods without requiring complex router training.
5.3 Ablation Studies and Sensitivity Analysis
We conduct comprehensive ablation studies to validate ABC’s design choices and analyze sensitivity to key deployment parameters:
Impact of Parallelization
§5.1.1 and Figure 8 (in Appendix E.1) analyze ABC under different parallelization scenarios. With full parallelization ($\rho=1$), ABC achieves optimal accuracy-FLOPs trade-offs, consistently outperforming single models. Even under sequential execution ($\rho=0$), ABC maintains substantial advantages when relative costs are sufficiently disparate (e.g., $\gamma\leq\frac{1}{50}$), demonstrating robustness to deployment constraints.
Relative Cost Sensitivity
§5.1.2 and Figure 3 examine how ABC depends on the relative cost $\gamma$ between cascade tiers. When models have similar costs ($\gamma\geq\frac{1}{5}$), some parallelization may be required for cost savings. However, when smaller models are at least 50$\times$ cheaper ($\gamma\leq\frac{1}{50}$), ABC provides meaningful savings even with sequential execution, validating our theoretical analysis.
Threshold Estimation Robustness
Appendix B and Figure 6 demonstrate that agreement threshold $\theta$ estimation is stable across different model accuracy levels (37.6% to 86.0%) and converges with minimal validation data. Using only 100 samples yields threshold estimates that remain stable when evaluated on 10$\times$ more data, confirming the effectiveness of the calibration procedure.
Safe Deferral Rule Existence
Appendix C and Figure 7 characterize selection rates under different error tolerances (1%, 3%, 5%) across model accuracies and computational budgets. Higher-accuracy models achieve selection rates up to 60% even with strict 1% error tolerance, empirically validating the practical existence of safe deferral rules predicted by our theory.
Cost Breakdown Analysis
Tables 4 & 5 (Appendix E.2) provide detailed analysis across cascade tiers, showing that most samples (52 to 93%) are processed at cheaper early tiers. This validates ABC’s ability to concentrate expensive computation on truly difficult samples while handling the majority of requests efficiently.
Cascade Configuration Effects
Figure 8 explores different cascade lengths (2-4 levels) and ensemble sizes (2-5 models per tier). The results show diminishing returns for larger ensembles, with 2-3 models per tier typically providing optimal accuracy-cost tradeoffs.
These ablations collectively demonstrate that ABC’s core design principles are empirically sound, with performance gracefully degrading under suboptimal deployment conditions while maintaining substantial benefits when key assumptions (e.g., large relative cost disparities) hold.
Takeaway #5: Ablation studies confirm that ABC’s benefits stem from the synergy of parallelization and cost disparities. Safe deferral rules exist across diverse settings, and threshold estimation requires minimal validation data ($\approx$100 samples) while remaining robust across model accuracies.